OpenStudy (anonymous):

Find the x-intercept(s): 4x^2-3y^2+2xy-12=0

OpenStudy (anonymous):

Do I just set the y=0 and solve from there?

OpenStudy (anonymous):

I got \[\pm \sqrt{3}\]

OpenStudy (anonymous):

Yes, just set y=0

OpenStudy (anonymous):

\[4x^2-3y^2+2xy-12=0\]

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):

yes thats correct method

OpenStudy (anonymous):

So your answer is correct

OpenStudy (cwrw238):

4x^2 = 12 z^2 = 3

OpenStudy (cwrw238):

* x^2 lol

OpenStudy (anonymous):

Thanks.